OpenStudy (jayslays):

Prove the Pythagorean Theorem using similar triangles. The Pythagorean Theorem states that in a right triangle, the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse. Be sure to create and name the appropriate geometric figures.

OpenStudy (leenathan):

gtg for the day now bye dan

OpenStudy (518nad):

later

OpenStudy (518nad):

hello @jayslays have u tried anything so far

OpenStudy (518nad):



OpenStudy (jayslays):

I want to have the hypotenuse equal 625, which is 25 squared. I'm just trying to find numbers that, squared, equal 625. I'm failing miserably. This is giving me a massive headache.

OpenStudy (518nad):

ah dont worry about it

OpenStudy (518nad):

u can easily find many number that work

OpenStudy (518nad):

625 = x^2+y^2 625 = 1 +624 so x^2=1 y^2=624 so x=sqrt1 y=sqrt624

OpenStudy (518nad):

this is one way to get 2 number that can square to 625 u can keep doing this pick and 2 positive number that add to 625

OpenStudy (518nad):

pick any*

OpenStudy (jayslays):

That makes sense. Thank you so much!

OpenStudy (518nad):

sure thing